# Generation of large-scale magnetic fields upstream of gamma-ray burst afterglow shocks

### v1

@@ 4.2 Asymptotic Behavior
While Equations (7)-(4.1) nicely capture the self-similarity of our system, we cannot use these equations to extrapolate our simulation data to arbitrary regimes due to the unknown functional forms of $\mathcal{F}$ and $\mathcal{G}$. That said, if these functions are weakly dependent on their arguments at late times (i.e., near $t=t_{f}$), we would expect the magnetic energy density and the transverse magnetic wavenumber at the end of pair injection to scale with the pre-factors in Equations (7)-(4.1): $\sigma_{f}$ (i.e., $\sigma$ at the end of pair injection) would scale with $\alpha^{1/2}\,\left(t_{f}\omega_{pi}\right)^{-1/2}$ and $k_{y}\right>_{f}c/\omega_{pi}$ would scale with $\alpha^{1/4}\,\left(t_{f}\omega_{pi}\right)^{-1/4}$. In this section, we measure the asymptotic scalings for the final magnetic energy and final transverse wavenumber, showing that the power-law dependencies on $t_{f}$ and $\alpha$ are not vastly different from those given by the self-similar scalings (7)-(4.1). In Section 5, we use these empirical scalings to extrapolate our simulation results to realistic $t_{f}$ and $\alpha$.
Figure 12 shows how $k_{y}\right>_{f}c/\omega_{pb}$ scales with $t_{f}\omega_{pb}$ and how $k_{y}\right>_{f}c/\omega_{pi}$ scales with $t_{f}\omega_{pi}$ for $\alpha=\{1,2,4,8,16\}$; both panels show the same simulations, but with the units appropriately converted. We do not include the $t_{f}\omega_{pb}=5\times 10^{3}$ (or $t_{f}\omega_{pi}\sim 117/\sqrt{\alpha}$) simulations in these plots, since we find that they generally do not follow the trends of the higher (and more realistic) $t_{f}\omega_{pb}$ cases (i.e., the $t_{f}\omega_{pb}=5\times 10^{3}$ cases are not yet in the asymptotic regime); for reference, we include the $t_{f}\omega_{pb}=5\times 10^{3}$ cases in Figure 18 of Appendix C. With $t_{f}\omega_{pb}=5\times 10^{3}$ discarded, the $k_{y}\right>_{f}c/\omega_{pb}$ dependence on $t_{f}\omega_{pb}$ at fixed $\alpha$ is well-fit by a power law $\propto\left(t_{f}\omega_{pb}\right)^{n}$, where $n$ ranges from $-0.50$ (for $\alpha=1$) to $-0.74$ (for $\alpha=16$); for $\alpha=2$, we obtain
$k_{y}\right>_{f}(\alpha=2)\sim 0.10\left(\frac{\omega_{pb}}{c}\right)\left(\frac{t_{f}\omega_{pb}}{10^{4}}\right)^{-0.6}$ (15)

### v2

@@ 4.2 Asymptotic Behavior
While Equations (7)-(4.1) nicely capture the self-similarity of our system, we cannot use these equations to extrapolate our simulation data to arbitrary regimes due to the unknown functional forms of $\mathcal{F}$ and $\mathcal{G}$. That said, if these functions are weakly dependent on their arguments at late times (i.e., near $t=t_{\rm f}$), we would expect the magnetic energy density and the transverse magnetic wavenumber at the end of pair injection to scale with the pre-factors in Equations (7)-(4.1): $\sigma_{\rm f}$ (i.e., $\sigma$ at the end of pair injection) would scale with $\alpha^{1/2}\,\left(t_{\rm f}\omega_{\rm pi}\right)^{-1/2}$ and $k_{y}\right>_{\rm f}c/\omega_{\rm pi}$ would scale with $\alpha^{1/4}\,\left(t_{\rm f}\omega_{\rm pi}\right)^{-1/4}$. In this section, we measure the asymptotic scalings for the final magnetic energy and final transverse wavenumber, showing that the power-law dependencies on $t_{\rm f}$ and $\alpha$ are not vastly different from those given by the self-similar scalings (7)-(4.1). In Section 5, we use these empirical scalings to extrapolate our simulation results to realistic $t_{\rm f}$ and $\alpha$.
Figure 12 shows how $k_{y}\right>_{\rm f}c/\omega_{\rm pb}$ scales with $t_{\rm f}\omega_{\rm pb}$ and how $k_{y}\right>_{\rm f}c/\omega_{\rm pi}$ scales with $t_{\rm f}\omega_{\rm pi}$ for $\alpha=\{1,2,4,8,16\}$; both panels show the same simulations, but with the units appropriately converted. We do not include the $t_{\rm f}\omega_{\rm pb}=5\times 10^{3}$ (or $t_{\rm f}\omega_{\rm pi}\sim 117/\sqrt{\alpha}$) simulations in these plots, since we find that they generally do not follow the trends of the higher (and more realistic) $t_{\rm f}\omega_{\rm pb}$ cases (i.e., the $t_{\rm f}\omega_{\rm pb}=5\times 10^{3}$ cases are not yet in the asymptotic regime); for reference, we include the $t_{\rm f}\omega_{\rm pb}=5\times 10^{3}$ cases in Figure 21 of Appendix E. With $t_{\rm f}\omega_{\rm pb}=5\times 10^{3}$ discarded, the $k_{y}\right>_{\rm f}c/\omega_{\rm pb}$ dependence on $t_{\rm f}\omega_{\rm pb}$ at fixed $\alpha$ is well-fit by a power law $\propto\left(t_{\rm f}\omega_{\rm pb}\right)^{n}$, where $n$ ranges from $-0.50$ (for $\alpha=1$) to $-0.74$ (for $\alpha=16$); for $\alpha=2$, we obtain
$k_{y}\right>_{\rm f}(\alpha=2)\sim 0.10\left(\frac{\omega_{\rm pb}}{c}\right)\left(\frac{t_{\rm f}\omega_{\rm pb}}{10^{4}}\right)^{-0.6}$ (15)